# Geometric Representation Condition Improves Equivariant Molecule Generation

### v1

@@ 4.2 Unconditional Molecule Generation
Main Results
We present the main results on the QM9 and DRUG datasets in Table 1. Below, we highlight the key findings: (i) Improvement over the base model (EDM): By leveraging geometric representations, GeoRCG significantly outperforms the base model, EDM, on both QM9 and DRUG datasets. Notably, on QM9, it increases stable molecules from $82\%$ to $93.9\%$ and validity from $91.9\%$ to $97.4\%$, while also improving molecule uniqueness. (ii) Superior performance compared to advanced methods: GeoRCG also surpasses included advanced models on the QM9 dataset. On the DRUG dataset, it outperforms models such as EDM-Bridge and GOAT, and gets a high score in validity. Although it falls short of achieving the best performance, we attribute this to the quality of the encoder’s representations, as we pre-trained it on the GEOM-DRUG dataset, which may lack diversity and sufficient size. Crucially, many structures in GEOM-DRUG lack the equilibrium conditions necessary for pre-training methods that enable effective learning of force fields (Zaidi et al., 2022; Feng et al., 2023). Additionally, GeoRCG adopts EDM (Hoogeboom et al., 2022) as the base molecule generator, which can be weaker compared to recent advanced models. Nevertheless, further improvement to GeoRCG can be achievabled by incorporating more advanced molecule generators or higher-quality pre-trained encoders.
Balancing Controllability
We proceed to investigate the “Balancing Controllablility”” feature of GeoRCG introduced in Section 3.2. To this end, we conducted a grid search by varying both $w$ and $\mathcal{T}$ on QM9 dataset, as depicted in Figure 4. The results indicate a clear trend: increasing $w$ and decreasing $\mathcal{T}$ improve validity and stability at the expense of uniqueness, allowing for fine-grained, flexible control over molecule generation. At its best, this approach achieves a molecule stability of 93.9% and a validity of 97.42%, approaching the dataset’s upper bound, with a trade-off in lower uniqueness&validity of 86.82%.
Additional Comparison with 2D&3D methods
We further compare with recent 2D&3D methods, including MiDi (Vignac et al., 2023) and LDM-3DG (You et al., 2023), which jointly learn and generate both 2D bond information and 3D geometry. However, these models rely on external tools like RDKit or OpenBabel (O’Boyle et al., 2011) for bond computation at the input stage, which enables them to implicitly leverage domain knowledge of such tools. In contrast, purely 3D methods like GeoRCG and those in our primary comparison utilize a coarse look-up table for bond estimation, which, while reflecting 3D learning, results in less accurate bond calculation, potentially biasing the comparison. Therefore, to provide a reference, we also report the performance of our models combined with the same external tools (e.g., OpenBabel) for precise bond calculations in our generated 3D conformations. Furthermore, as GeoRCG essentially captures 3D geometric distributions, we place more emphasis on 3D metrics that directly evaluate 3D learning capabilities, including BondLengthW1 and BondAngleW1 proposed by Vignac et al. (2023) and detailed in Appendix C. The results in Table 2 demonstrate that GeoRCG not only significantly outperforms MiDi and LDM-3DG on 3D metrics, highlighting the advantages of using a pure 3D model for learning 3D structures, but also further enhances EDM’s performance, which has already shown considerable promise in 3D learning.
Data $\sim$0.1 $\sim$0 98.7 99.8 98.9 99.9
LDM-3DG 3.56 0.2 94.03 99.38 94.89 97.03
EDM 0.44 0.1 90.7 99.2 91.7 98.5
EDM + OBabel 0.44 0.1 97.9 99.8 99.0 98.5
4.3 Conditional Molecule Generation
@@
Figure 5: Conditionally generated molecules on property $\alpha$ using GeoRCG. The black number indicates the specified property value (condition), the green number represents the evaluated property value (computed by the classifier) for the generated molecule conformer.
@@ 4.4 Faster Generation with Fewer Steps
With geometric representation condition, it is reasonable to expect that fewer discretization steps of the reverse diffusion SDE (Song et al., 2021) would still yield competitive results. Therefore, we reduce the number of diffusion steps and evaluate the model’s performance. The results are presented in Table 4. As demonstrated, with the geometric representation condition, GeoRCG consistently outperforms other approaches across all step numbers. Notably, with approximately 100 steps, the performance of our method nearly converges to the optimal performance observed with 1000 steps, which already surpasses all other methods across all step numbers. This directly reflects the faster generation speed, as the time required for the first-stage representation generation is minimal and can be considered negligible.
Methods Metrics # Steps Atom Sta (%) $\uparrow$ Mol Sta (%) $\uparrow$ Valid (%) $\uparrow$ Valid & Unique (%) $\uparrow$
Data - 99 95.2 97.7 97.7
Best baseline in Table 1 1000 99.08 90.87 95.31 92.96
EquiFM 200 98.9(0.1) 88.3(0.3) 94.7(0.4) 93.5(0.3)
GOAT 90 98.4 84.1 90.9 89.99
EDM 50 97.0(0.1) 66.4(0.2) - -
EDM-Bridge 50 97.3(0.1) 69.2(0.2) - -
GeoBFN 50 98.28(0.1) 85.11(0.5) 92.27(0.4) 90.72(0.3)
GeoRCG 50 98.75(0.05) 1.80% 89.08(0.52) 34.16% 95.05(0.33) 91.32(0.37)
EDM 100 97.3(0.1) 69.8(0.2) - -
EDM-Bridge 100 97.9(0.1) 72.3(0.2) - -
GeoBFN 100 98.64(0.1) 87.21(0.3) 93.03(0.3) 91.53(0.3)
GeoRCG 100 99.08(0.03) 1.83% 91.85(0.34) 31.59% 96.49(0.27) 92.07(0.35)
EDM 500 98.5(0.1) 81.2(0.1) - -
EDM-Bridge 500 98.7(0.1) 83.7(0.1) - -
GeoBFN 500 98.78(0.8) 88.42(0.2) 93.35(0.2) 91.78(0.2)
GeoRCG 500 99.09(0.01) 0.60% 91.89(0.24) 13.17% 96.57(0.12) 92.08(0.36)
EDM 1000 98.7 82 91.9 90.7
EDM-Bridge 1000 98.8 84.6 92 90.7
GeoBFN 1000 99.08(0.06) 90.87(0.2) 95.31(0.1) 92.96(0.1)
GeoRCG 1000 99.12(0.03) 0.43% 92.32(0.06) 12.59% 96.52(0.2) 5.03% 92.45(0.2) 1.93%
5 Conclusions
@@
In this work, we present GeoRCG, an effective framework for improving the generation quality of arbitrary molecule generators by incorporating geometric representation conditions. We use EDM (Hoogeboom et al., 2022) as the base molecule generator and demonstrate the effectiveness of our framework through extensive molecular generation experiments. Notably, in conditional generation tasks, GeoRCG achieves a 31% performance boost compared to recent state-of-the-art models. Additionally, the representation guidance enables significantly faster sampling with 10x fewer diffusion steps while maintaining near-optimal performance. Beyond these empirical improvements, we provide theoretical characterizations of general representation-conditioned generative models, which address a key gap in the existing literature (Li et al., 2023).
One limitation of our framework is that its generation quality could depend on the representation quality. For instance, on the GEOM-DRUG dataset, where the encoder was less thoroughly pre-trained, the improvements were less pronounced, and GeoRCG did not surpass SOTA methods. Future work could focus on improving the effectiveness of pre-trained models or exploring enhanced representation regularization techniques. Furthermore, while we employ EDM (Hoogeboom et al., 2022) as the base molecule generator, our framework is general and can be applied to any molecular generative model. Integrating this framework with more advanced SOTA models offers a promising direction for future exploration and performance enhancement.
@@ Appendix B Algorithms
Input: Molecule dataset $\mathcal{D}^{\text{mol}}_{\text{train}}\subset\bigcup_{N=1}^{+\infty}\left(\mathbb{R}^{N\times 3}\times\mathbb{R}^{N\times d}\right)$, pre-trained geometric encoder $E$, initial representation generator $p_{\varphi_{0}}(r)$ and molecule generator $p_{\theta_{0}}(\mathcal{M}|r)$, classifier-free guidance coefficient $w$, temperature $\mathcal{T}$.
Output: Trained representation generator $p_{\varphi}(r)$ and molecule generator $p_{\theta}(\mathcal{M}|r)$, molecule samples from $p_{\varphi,\theta}(\mathcal{M})$.
1:Parallel Training
2:Pre-process to obtain the representation dataset $\mathcal{D}_{\text{train}}^{\text{rep}}=\{\!\!\{(E(\mathcal{M}),N(\mathcal{M}))|\mathcal{M}\in\mathcal{D}^{\text{mol}}_{\text{train}}\}\!\!\}$ and the mol-rep dataset $\mathcal{D}_{\text{train}}^{\text{mol-rep}}=\{\!\!\{(E(\mathcal{M}),\mathcal{M})|\mathcal{M}\in\mathcal{D}^{\text{mol}}_{\text{train}}\}\!\!\}$.
3:(Parallelly) Train the representation generator with $\mathcal{D}^{\text{rep}}_{\text{train}}$ with loss $\mathcal{L}_{\text{rep}}$ defined in Equation 1; Train the molecule generator with $\mathcal{D}^{\text{mol-rep}}_{\text{train}}$ with loss $\mathcal{L}_{\text{mol}}$ defined in Equation 2, along with representation perturbation introduced in Appendix B.
1:Sequential Sampling
2:Sample a representation $r_{*}\sim p_{\varphi}(r)$ with temperature control $\mathcal{T}$.
3:Sample a molecule $\mathcal{M}_{*}\sim p_{\theta}(\mathcal{M}|r_{*})$, conditionally with classifier-free guidance $w$.
4:Return: Trained representation generator $p_{\varphi}(r)$, molecule generator $p_{\theta}(\mathcal{M}|r)$, and generated molecule sample.
@@ Appendix C Experiment Details
    When comparing with 2D & 3D models, we evaluate two 3D metrics introduced by MiDi (Vignac et al., 2023), which directly assess the geometry learning ability:
@@
  • Representation Generator. We use the same architecture for the representation generator as the MLP-based diffusion model proposed in Li et al. (2023). We use 18 blocks of residual MLP layers with 1536 hidden dimensions, 1000 diffusion steps, and a linear noise schedule for $\beta_{t}$. The representation generator is trained for 2000 epochs with a batch size of 128 for both the QM9 and DRUG datasets. Training on QM9 takes approximately 2.5 days on a single Nvidia 4090, while training on DRUG takes around 4 days on a single Nvidia A800. Training time can indeed be further reduced, as the model shows minimal progress after approximately half of the reported time.
@@
  • Molecule Generator. We adopt EDM (Hoogeboom et al., 2022) as the base molecule generator, using the same EGNN (Satorras et al., 2021) architecture, with the exception of the conditioning module. Specifically, we incorporate a cross-attention (Vaswani, 2017) module for better conditioning on representations, placing it between every block of the EGNN to enhance regularization and increase expressiveness. For the EGNN hyperparameters, we use 9 layers with 256 hidden dimensions for QM9 and 4 layers with 256 hidden dimensions for DRUG. The number of diffusion steps is set to 1000 (except for cases in Table 4 that generate molecules with fewer steps), and we employ the polynomial scheduler for $\alpha_{t}^{(\mathcal{M})}$. Notably, all model hyperparameters are identical to those in EDM for fair comparison. During training, we use a batch size of 128 and 3000 epochs on QM9, and a batch size of 64 and 20 epochs on DRUG. The representation perturbation values are set to $\sigma_{\text{rep}}=0.3$ on QM9 and $\sigma_{\text{rep}}=0.5$ on DRUG. Training takes approximately 6 days on QM9 using a single Nvidia 4090, and around 10 days on DRUG using two Nvidia A800 GPUs.
@@
  • Sampling Details. During representation sampling, we use the predictor-corrector method from Song et al. (2021) for VP SDE to achieve better equilibrium distributions, which facilitates low-temperature sampling. For molecule sampling, the process remains the same as in EDM, with the addition of classifier-free guidance. For the unconditional generation results in Table 1, we use $(w,\mathcal{T})=(1.0,1.0)$ in the QM9 dataset, except for the results in Figure 4, and $(w,\mathcal{T})=(0.0,0.5)$ in the GEOM-DRUG dataset (note that $w=0.0$ indicates that we do not use classifier-free guidance, not that the representation condition itself is not applied). Further tuning of these two hyperparameters may lead to improved results.
@@
We evaluate the performance of LDM-3DG (You et al., 2023), an Auto-Encoder-based method that also leverages the compactness of the representation space to achieve good performance.

### v2

@@ 4.2 Unconditional Molecule Generation
We present the main results on the QM9 and GEOM-DRUG datasets in Table 1. Below, we highlight the key findings: (i) Improvement over the base model (EDM): By leveraging geometric representations, GeoRCG significantly outperforms the base model, EDM, on both QM9 and GEOM-DRUG datasets. Notably, on QM9, it increases stable molecules from $82\%$ to $93.9\%$ and validity from $91.9\%$ to $97.4\%$, while also improving molecule uniqueness. (ii) Superior performance compared to advanced methods: GeoRCG also surpasses included advanced models on the QM9 dataset. On the GEOM-DRUG dataset, it outperforms models such as EDM-Bridge and GOAT, and gets a high score in validity. Although it falls short of achieving the best performance, we attribute this to the relatively limited capabilities of EDM. To address this, we replace EDM with the recent SOTA flow-matching based model, SemlaFlow (Irwin et al., 2024), as the base model on the GEOM-DRUG dataset, as shown in Table 3. As demonstrated, GeoRCG consistently enhances SemlaFlow’s SOTA performance across all metrics on the GEOM-DRUG dataset.
Table 3: GeoRCG with SemlaFlow as the backbone. Molecule stability and validity are reported as percentages, while energy and strain energy are expressed in kcal·mol-1. Results marked with were reproduced in our own experiments. All results are calculated based on 5k randomly sampled molecules, averaged over five runs.
Methods Atom Stab $\uparrow$ Mol Stab $\uparrow$ Valid $\uparrow$ Energy $\downarrow$ Strain $\downarrow$
MiDi 99.8 91.6 77.8 - -
EQGAT-diff 99.8(0.01) 93.4(0.21) 94.6(0.24) 148.8(0.9) 140.2(0.7)
SemlaFlow 99.9(0.00) 97.6(0.2) 94.0(0.2) 93.72(1.06) 52.98(0.87)
GeoRCG (Semla) 99.9(0.00) 97.7(0.1) 95.1(0.22) 87.86(0.33) 47.07(0.81)
Table 4: Quality of generated molecules on GEOM-DRUG with fewer diffusion steps.
# Steps Atom Sta (%) $\uparrow$ Valid (%) $\uparrow$
GeoBFN GeoRCG GeoBFN GeoRCG
50 75.11 81.44(0.10) 91.66 95.70(0.70)
100 78.89 83.02(0.06) 93.05 96.30(0.70)
500 81.39 84.03(0.37) 93.47 97.57(0.90)
1000 85.6 84.3(0.12) 92.08 98.5(0.12)
In Appendix D, we present additional experiments on QM9, demonstrating that: 1) GeoRCG effectively balances uniqueness and validity through the controllability feature introduced in Section 3.2, and 2) GeoRCG enhances distribution-level geometric metrics, such as BondAngleW1, which calculates the 1-Wasserstein distance between the atom-centered angle distribution of the generated molecules and that of the training dataset, which underscore GeoRCG’s improved geometric learning capabilities.
4.3 Conditional Molecule Generation
@@
Methods Metrics # Steps Atom Sta (%) $\uparrow$ Mol Sta (%) $\uparrow$ Valid (%) $\uparrow$
Data - 99 95.2 97.7
EquiFM 200 98.9(0.1) 88.3(0.3) 94.7(0.4)
GOAT 90 98.4 84.1 90.9
EDM 50 97.0(0.1) 66.4(0.2) -
EDM-Bridge 50 97.3(0.1) 69.2(0.2) -
GeoBFN 50 98.28(0.1) 85.11(0.5) 92.27(0.4)
GeoRCG 50 98.75(0.05) 1.80% 89.08(0.52) 34.16% 95.05(0.33)
EDM 100 97.3(0.1) 69.8(0.2) -
EDM-Bridge 100 97.9(0.1) 72.3(0.2) -
GeoBFN 100 98.64(0.1) 87.21(0.3) 93.03(0.3)
GeoRCG 100 99.08(0.03) 1.83% 91.85(0.34) 31.59% 96.49(0.27)
EDM 500 98.5(0.1) 81.2(0.1) -
EDM-Bridge 500 98.7(0.1) 83.7(0.1) -
GeoBFN 500 98.78(0.8) 88.42(0.2) 93.35(0.2)
GeoRCG 500 99.09(0.01) 0.60% 91.89(0.24) 13.17% 96.57(0.12)
EDM 1000 98.7 82 91.9
EDM-Bridge 1000 98.8 84.6 92
GeoBFN 1000 99.08(0.06) 90.87(0.2) 95.31(0.1)
GeoRCG 1000 99.12(0.03) 0.43% 92.32(0.06) 12.59% 96.52(0.2) 5.03%
Figure 3: Conditionally generated molecules on property $\alpha$ using GeoRCG. The black number indicates the condition value, the green number represents the oracle property value for the generated molecule conformer.
@@ 4.4 Faster Generation with Fewer Steps
With geometric representation condition, it is reasonable to expect that fewer discretization steps of the reverse diffusion SDE (Song et al., 2021) would still yield competitive results. Therefore, we reduce the number of diffusion steps and evaluate the model’s performance. The results are presented in Table 5 and Table 4. As demonstrated, with the geometric representation condition, GeoRCG consistently outperforms other approaches across almost all step numbers. Notably, in Table 5, with approximately 100 steps, the performance of our method nearly converges to the optimal performance observed with 1000 steps, which already surpasses all other methods across all step numbers. This directly reflects the faster generation speed, as the time required for the first-stage representation generation is minimal and can be considered negligible.
5 Conclusions and Limitations
@@
Conclusions. In this work, we present GeoRCG, an effective framework for improving the generation quality of arbitrary molecule generators by incorporating geometric representation conditions. We use EDM (Hoogeboom et al., 2022) and SemlaFlow (Irwin et al., 2024) as base molecule generators and demonstrate the effectiveness of our framework through extensive molecular generation experiments. Notably, in conditional generation tasks, GeoRCG achieves a 31% performance boost compared to recent SOTA models. Additionally, the representation guidance enables significantly faster sampling with 10x fewer diffusion steps while maintaining near-optimal performance. Beyond these empirical improvements, we provide theoretical characterizations of representation-conditioned generative models, which address a key gap in the existing literature (Li et al., 2023).
Limitations. A key limitation of our framework is that generation quality may depend heavily on the quality of representations. For instance, in the GEOM-DRUG dataset, an insufficiently pre-trained encoder may produce less meaningful representations. As a result, the benefits of low-temperature sampling and classifier-free guidance in enhancing generation quality and controllability may be less pronounced. Future work could investigate more effective pre-training strategies beyond standard denoising or enhanced representation regularization techniques to mitigate this issue.
Impact Statement
This paper presents work whose goal is to advance the field of molecule generation. There are many potential societal consequences of molecule generation improvement, such as accelerating drug discovery, and developing new material. None of them we feel need to be specifically highlighted here for potential risk.
@@ Appendix B Algorithms
  Input: Molecule dataset $\mathcal{D}^{\text{mol}}_{\text{train}}\subset\bigcup_{N=1}^{+\infty}\left(\mathbb{R}^{N\times 3}\times\mathbb{R}^{N\times d}\right)$, pre-trained geometric encoder $E$, initial representation generator $p_{\varphi_{0}}(r)$, molecule generator $p_{\theta_{0}}(\mathcal{M}|r)$, classifier-free guidance coefficient $w$, temperature $\mathcal{T}$.
  Output: Trained representation generator $p_{\varphi}(r)$, molecule generator $p_{\theta}(\mathcal{M}|r)$, and molecule samples from $p_{\varphi,\theta}(\mathcal{M})$.
  Parallel Training:
  Pre-process to obtain:
      - The representation dataset $\mathcal{D}_{\text{train}}^{\text{rep}}=\{\!\!\{(E(\mathcal{M}),N(\mathcal{M}))|\mathcal{M}\in\mathcal{D}^{\text{mol}}_{\text{train}}\}\!\!\}$
      - The mol-rep dataset $\mathcal{D}_{\text{train}}^{\text{mol-rep}}=\{\!\!\{(E(\mathcal{M}),\mathcal{M})|\mathcal{M}\in\mathcal{D}^{\text{mol}}_{\text{train}}\}\!\!\}$
  Train the representation generator with $\mathcal{D}^{\text{rep}}_{\text{train}}$ using loss $\mathcal{L}_{\text{rep}}$.
  Train the molecule generator with $\mathcal{D}^{\text{mol-rep}}_{\text{train}}$ using loss $\mathcal{L}_{\text{mol}}$ and representation perturbation.
  Sequential Sampling:
  Sample a representation $r_{*}\sim p_{\varphi}(r)$ with temperature control $\mathcal{T}$.
  Sample a molecule $\mathcal{M}_{*}\sim p_{\theta}(\mathcal{M}|r_{*})$ conditionally, with classifier-free guidance $w$.
  Return: Trained representation generator $p_{\varphi}(r)$, molecule generator $p_{\theta}(\mathcal{M}|r)$, and generated molecule sample.
@@ Appendix C Experiment Details
    • Energy: The energy $U(x)$ of a conformation $x$. Lower energy values typically correspond to more stable and physically plausible conformations that are closer to what would be observed in nature. Following (Irwin et al., 2024), the energy is calculated using the MMFF94 force field within RDKit, a commonly used molecular modeling framework.
    • Strain: A measure of how distorted a generated conformation $x$ is compared to its relaxed (optimized) state $\tilde{x}$. The relaxed conformation $\tilde{x}$ is obtained by applying energy minimization using the MMFF94 force field, where the molecular structure is iteratively adjusted to reduce its energy until reaching a local minimum. Mathematically, the strain energy is defined as $U(x)-U(\tilde{x})$ , where $U(x)$ is the energy of the generated conformation and $U(\tilde{x})$ is the energy of the relaxed conformation. Lower strain energy values imply that the generated conformations are closer to being physically accurate and require minimal correction during optimization.
@@
    When comparing with 2D & 3D models in Table 7, we evaluate two 3D metrics introduced by MiDi (Vignac et al., 2023), which directly assess the geometry learning ability:
@@
  • Representation Generator. We use the same architecture for the representation generator as the MLP-based diffusion model proposed in Li et al. (2023). We use 18 blocks of residual MLP layers with 1536 hidden dimensions, 1000 diffusion steps, and a linear noise schedule for $\beta_{t}$. The representation generator is trained for 2000 epochs with a batch size of 128 for both the QM9 and GEOM-DRUG datasets. Training on QM9 takes approximately 2.5 days on a single Nvidia 4090, while training on GEOM-DRUG takes around 4 days on a single Nvidia A800. Training time can indeed be further reduced, as the model shows minimal progress after approximately half of the reported time.
@@
  • Molecule Generator. We adopt EDM (Hoogeboom et al., 2022) as the base molecule generator, using the same EGNN (Satorras et al., 2021) architecture, with the exception of the conditioning module. Specifically, we incorporate a cross-attention (Vaswani, 2017) module for better conditioning on representations, placing it between every block of the EGNN to enhance regularization and increase expressiveness. For the EGNN hyperparameters, we use 9 layers with 256 hidden dimensions for QM9 and 4 layers with 256 hidden dimensions for GEOM-DRUG. The number of diffusion steps is set to 1000 (except for cases in Table 5 that generate molecules with fewer steps), and we employ the polynomial scheduler for $\alpha_{t}^{(\mathcal{M})}$. Notably, all model hyperparameters are identical to those in EDM for fair comparison. During training, we use a batch size of 128 and 3000 epochs on QM9, and a batch size of 64 and 20 epochs on GEOM-DRUG. The representation perturbation values are set to $\sigma_{\text{rep}}=0.3$ on QM9 and $\sigma_{\text{rep}}=0.5$ on GEOM-DRUG. Training takes approximately 6 days on QM9 using a single Nvidia 4090, and around 10 days on GEOM-DRUG using two Nvidia A800 GPUs.
    For GeoRCG (SemlaFlow), we similary adopt the cross-attention module for conditioning on representations and place it between every block of Semla, which is the backbone model of SemlaFlow. During training, we use the batch cost 2048 and trains for 300 epochs. We set the $\sigma_{\text{rep}}=0.3$. Notably, we introduce an auxiliary loss, referred to as the representation loss, which is computed as the MSE between the representations (calculated by the pre-trained encoder) of the ground-truth molecule and the denoised molecule. This loss is conceptually similar to the denoising loss but operates in the representation space rather than the raw molecular space. To balance its contribution, we scale the representation loss by a coefficient of 0.5.
@@
  • Sampling Details. During representation sampling, we use the predictor-corrector method from Song et al. (2021) for VP SDE to achieve better equilibrium distributions, which facilitates low-temperature sampling. For molecule sampling, the process remains the same as in EDM, with the addition of classifier-free guidance. For the unconditional generation results in Table 1, we use $(w,\mathcal{T})=(1.0,1.0)$ in the QM9 dataset, except for the results in Figure 5, and $(w,\mathcal{T})=(0.0,0.5)$ in the GEOM-DRUG dataset. For GeoRCG (SemlaFlow), we set $(w,\mathcal{T})=(-0.9,1.0)$ (note that $-1.0<w<0.0$ indicates that the representation guidance is present but subtle, exerting a small yet meaningful influence on the model’s behavior). Further tuning of these two hyperparameters may lead to improved results.
@@
We evaluate the performance of LDM-3DG (You et al., 2023) in Table 7 and Table 2, an Auto-Encoder-based method that also leverages the compactness of the representation space to achieve good performance.